# Multi-Cache Enhanced Prototype Learning for Test-Time Generalization of Vision-Language Models

### v1

@@ Abstract
In zero-shot setting, test-time adaptation adjusts pre-trained models using unlabeled data from the test phase to enhance performance on unknown test distributions. Existing cache-enhanced TTA methods rely on a low-entropy criterion to select samples for prototype construction, assuming intra-class compactness. However, low-entropy samples may be unreliable under distribution shifts, and the resulting prototypes may not ensure compact intra-class distributions. This study identifies a positive correlation between cache-enhanced performance and intra-class compactness. Based on this observation, we propose a Multi-Cache enhanced Prototype-based Test-Time Adaptation (MCP) featuring three caches: an entropy cache for initializing prototype representations with low-entropy samples, an align cache for integrating visual and textual information to achieve compact intra-class distributions, and a negative cache for prediction calibration using high-entropy samples. We further developed MCP++, a framework incorporating cross-modal prototype alignment and residual learning, introducing prototype residual fine-tuning. Comparative and ablation experiments across 15 downstream tasks demonstrate that the proposed method and framework achieve state-of-the-art generalization performance.
@@ 3.2 Multi-Cache Enhanced Prototype
where $H_{\max}^{\hat{y}}$ denotes the highest entropy value among samples stored in the align cache for category $\hat{y}$. If this condition is met, we further calculate the Euclidean distance between the feature embedding of the test sample $f_{\text{test}}$ and the corresponding category prototype center $\mu_{\hat{y}}$, denoted as $d(f_{\text{test}},\mu_{\hat{y}}),$ and compare it with the distance of the highest-entropy cached sample $f_{\max}$ to the prototype center, $d(f_{\max},\mu_{\hat{y}})$. Only if
$d(f_{\text{test}},\mu_{\hat{y}})<d(f_{\max},\mu_{\hat{y}})$
@@ E.4 Assumptions
[…]
@@
[…]
@@ E.6 Propositions
[…]
@@
$(\displaystyle\mathcal{E}(f)\;\leq\;\;\mathcal{O}\Bigl{(}$$\bigl{(}\tfrac{1}{k_{t}+k_{a}}\bigr{)}^{\tfrac{1}{4}}\;+\;\sum_{i=1}^{k_{t}}w_{ti}\,\bigl{(}\tfrac{k_{t}}{c_{t}\,n_{t}}\bigr{)}^{\tfrac{1}{d}}$ (17)
[…]
@@ E.7 Proof of Proposition 2 Improvement
Let the filtering rule select samples satisfying event $A$ (Here, entropy is less than $\leq\tau$), which is statistically independent of the alignment condition =$\mathcal{R}=\{x:\|g(x)-\mu_{c}\|\leq d_{0}\}$. Then:
@@
    $(\displaystyle\mathcal{E}(f)\;\leq\;\;\mathcal{O}\Biggl{(}$$\underbrace{\biggl{(}\frac{1}{k_{t}+k_{a}}\biggr{)}^{\!\tfrac{1}{4}}}_{\text{\rm Term' 1}}$
@@
    $\underbrace{\text{Term' 1 + Term' 2}}_{\text{Proposition 2}}\;<\;\underbrace{\text{Term 1 + Term 2}}_{\text{Proposition 1}}.$

### v2

@@ Abstract
In zero-shot setting, test-time adaptation adjusts pre-trained models using unlabeled data from the test phase to enhance performance on unknown test distributions. Existing cache-enhanced TTA methods rely on a low-entropy criterion to select samples for prototype construction, assuming intra-class compactness. However, low-entropy samples may be unreliable under distribution shifts, and the resulting prototypes may not ensure compact intra-class distributions. This study identifies a positive correlation between cache-enhanced performance and intra-class compactness. Based on this observation, we propose a Multi-Cache enhanced Prototype-based Test-Time Adaptation (MCP) featuring three caches: an entropy cache for initializing prototype representations with low-entropy samples, an align cache for integrating visual and textual information to achieve compact intra-class distributions, and a negative cache for prediction calibration using high-entropy samples. We further developed MCP++, a framework incorporating cross-modal prototype alignment and residual learning, introducing prototype residual fine-tuning. Comparative and ablation experiments across 15 downstream tasks demonstrate that the proposed method and framework achieve state-of-the-art generalization performance.Project Page available at: https://zhaihaotian.github.io/MCP-ICCV25/
@@ 3.2 Multi-Cache Enhanced Prototype
[…]
@@ E.4 Assumptions
$\left\{\begin{aligned} &\lambda\bigl{[}p_{t}(x)\cap\mathcal{B}(x_{0},r)\bigr{]}\;\geq\;c_{t}\,\lambda\bigl{[}\mathcal{B}(x_{0},r)\bigr{]},\\[3.0pt] &\lambda\bigl{[}p_{a}(x_{0})\cap\mathcal{B}(x_{0},r)\bigr{]}\;\geq\;c_{a}\,\lambda\bigl{[}\mathcal{B}(x_{0},r)\bigr{]},\\[3.0pt] &m\;<\;\tfrac{dp_{t}(x)}{d\lambda}\;<\;M,\quad m\;<\;\tfrac{dp_{a}(x)}{d\lambda}\;<\;M.\end{aligned}\right$.
@@
$\begin{split}\lambda\Bigl{[}p_{a}(x)\cap\mathcal{B}(x_{0},r)\Bigr{]}&=\int_{\mathcal{B}(x_{0},r)}p_{a}(x)\,dx\\ &=\frac{1}{P(\|g(x)-\mu_{c}\|\leq d_{0})}\int_{\mathcal{B}(x_{0},r)}p_{t}(x)\,dx.\end{split}$ (15)
@@ E.6 Propositions
$\mathcal{E}(f)\;\leq\;\mathcal{O}\!\Bigl{(}\bigl{(}\tfrac{1}{k_{t}}\bigr{)}^{\tfrac{1}{4}}\;+\;\bigl{(}\tfrac{k_{t}}{c_{t}\,n_{t}}\bigr{)}^{\tfrac{1}{d}}\Bigr{)}^{\,1+\beta}.$ (16)
@@
$\mathcal{E}(f)\;\leq\;\;\mathcal{O}\Bigl{(}$ $\bigl{(}\tfrac{1}{k_{t}+k_{a}}\bigr{)}^{\tfrac{1}{4}}\;+\;\sum_{i=1}^{k_{t}}w_{ti}\,\bigl{(}\tfrac{k_{t}}{c_{t}\,n_{t}}\bigr{)}^{\tfrac{1}{d}}$ (17)
$+\;\sum_{i=1}^{k_{a}}w_{ai}\,\bigl{(}\tfrac{k_{a}}{c_{a}\,n_{a}}\bigr{)}^{\tfrac{1}{d}}\Bigr{)}^{\,1+\beta}.$
@@ E.7 Proof of Proposition 2 Improvement
Let the filtering rule select samples satisfying event $A$ (Here, entropy is less than $\leq\tau$), which is statistically independent of the alignment condition $\mathcal{R}=\{x:\|g(x)-\mu_{c}\|\leq d_{0}\}$. Then:
@@
    $\mathcal{E}(f)\;\leq\;\;\mathcal{O}\Biggl{(}$ $\underbrace{\biggl{(}\frac{1}{k_{t}+k_{a}}\biggr{)}^{\!\tfrac{1}{4}}}_{\text{\rm Term' 1}}$
@@
[…]